# Frame In-N-Out: Unbounded Controllable Image-to-Video Generation

### v1

@@ Appendix C Additional Experimental Details
In the motion control architecture, one-pixel coordinate as motion images form is insensitive for the VAE encoder to effectively encode and decode; thus, similar to previous works [68, 59], we enlarge pixels to a square rectangular box and then execute a 2D dilation algorithm from deblurring domain to increase the perceptibility. The rectangular box border length is 6 pixels for a video with an original height of 384, which means that the size will be scaled with the exact height. Further, to accommodate sparse motion points (1-2 points per object) provided by the user side during the inference, we randomly drop points during the training. The number of tracking points is randomly dropped with the probability of 33% and 60% for each stage respectively. Different from optical flow-based motion control works [68, 59, 87], we do not need a padding frame on the last frame to align the total number of frames, where all of our motion conditioning frames represent a spatiotemporal pixel-aligned motion status.
Our inference pipeline can be found in Fig. 5. The model will start from pure noise with the size of the canvas region. The first frame will be expanded to the canvas size by the $\mathbf{B}_{\text{canvas}}$ setting for the top-left and bottom-right expansion pixel number. The motion is conditioned on the canvas size as a whole after converting from the raw pixel spatiotemporal coordinates. Similarly, the ID reference images will be scaled and padded to the same size as the canvas size. The denoised latent will be decoded by the pre-trained VAE decoder and then we cropped the pixel space videos to be the exact same shape as the first frame based on $\mathbf{B}_{\text{canvas}}$ setting.
@@
Since the resolution and the number of frames that each baseline model can generate are different in the quantitative comparisons. For a fair comparison, we resize the validation dataset to the original process resolution of each model. Then, we resize the generated videos to the uniform resolution of 256$\times$384 with the number of frames of 14 and 49 respectively for the Frame Out and Frame In table. This number of frames represents the minimum frame number across all models in that table. For the SkyReels-A2 [13], based on their official code implementation, we generate 81 frames and then crop the first 12 frames since the first 12 frames look highly distorted for an unfair comparison. We only compare the remaining 69 frames.
Works like ReVideo [38] and VideoAnyDoor [54] and other similar worsk [92, 6] are conditioned on the full video sequence as inputs, not just the first frame, which means that they are video-to-video generation instead of image-to-video generation we target at. Furthermore, these works are more focused on the insertion and editing in the middle instead of entering and leaving effects we pursue. Thus, we believe that this direction is not an ideal choice for Frame In and Frame Out baselines. An additional condition comparison table can be found in Tab. 5.
@@
FID [19] measures the distance between the distribution of generated and real frame features, extracted using an Inception network. It is sensitive to both image quality and diversity, and lower scores indicate better generation.
FVD [56] extends FID to the temporal domain, computing the Fréchet distance between distributions of video-level features extracted using a pre-trained I3D model. It captures spatiotemporal coherence in generated videos and is similarly lower-is-better.
LPIPS [84] quantifies the perceptual similarity between individual frames and their references using deep feature distances from a pre-trained network. Unlike pixel-wise metrics, LPIPS correlates well with human judgment of visual similarity.
The prompt we used for motion description for SkyReels-A2 [13] and Phantom [36] is: Please describe the video in 50 words. Describe the motion of the object clearly and in details, but in the natural and direct language. It is expected that an object will enter and/or exit the image. Describe how the character is moved in and exit, like from the top, left, right, bottom.
@@ Appendix D Ablation Study
Table 6: Ablation Study on Frame In.
@@
As shown in Tab. 6, we compare several different purposes-trained models under different settings. Due to the computation limitations, we train all models with 12K iterations of a batch size of 8. The training is done on stage 2 and all utilize the same pre-trained stage 1 weight (including the baseline). We compare all models on the Frame In evaluation, which is the most representative task for all conditions to be considered.
@@
Full Field Loss Influence. We also provide a model that is not trained with full field loss. This means that the ground-truth target latent provided only has the encoded information from the first frame region instead of the canvas region in the baseline. The region outside the first frame is padded with zero. It is worth noting that the comparison of FID, FVD, LPIPS, Relative DINO, and VLM are only the first frame region, which means that the padded zero is not included in the evaluation. As we can see from the third row of the Tab. 6, almost all metrics dropped. Here, trajectory error, video segmentation MAE, and relative DINO dropped evidently. We believe that the introduction of full field loss is significant to the final visual generated quality and motion consistency.
Absolute Position Encoding Influence. In our baseline model, we train the model with resized absolute position encoding by trilinear interpolation. Despite this method, another solution to embrace different resolution inputs for both training and inference could be to create a new position encoding each time, where the position encoding for the identity reference is also refreshed each time. However, as shown in the fourth row of Tab. 6, this will lead to a performance drop in all metrics, especially FID and relative DINO. We believe that reusing the learned position encoding by applying a simple trilinear interpolation is helpful to the versatile video resolution inputs, which maintain similar data distribution from the learned fixed position embeddings at the minimum cost.
@@ Appendix E Additional Visualization
Figure 7: More Generated Examples.
A more complete sequence of the generated videos and all conditions of the teaser can be found in Fig. 6. Further, we provide more of our generated video samples in Fig. 7. Some demo images or ID reference is chosen or cropped from Davis Dataset [42] and online images. We show multiple different kinds of combinations of Frame In and Frame Out. For the physical interaction, we mean the interaction caused by breaking new ID references to the object that already exists in the first frame. We set all the height and width to be a multiplier of 32 due to the VAE limitation. As shown in Fig. 6 and Fig. 7, the setting for the height and width of the canvas and first frame region can be versatile in aspect ratio and size while generating high fidelity and stable results. Further, we empirically found the generation is stable as long as the canvas height is less than 480 pixels and the canvas width is less than 720 pixels, which is the training resolution for our pre-trained base model, CogVideoX [77].
As shown in Fig. 8 and Fig. 9, we also present extra qualitative comparisons with the baseline methods. For the Frame In comparison, we can see that baselines like Phantom [36] and SkyReels-A2 [13] cannot understand the Frame In intention we want. The identity reference either already existed in the first frame (case 2 and 3) or never coming to the scene (case 1). Further, the first frame condition is not faithfully considered as the main condition. The scene they generate is similar in elements but different from objects, 3D position, and physical relationship. On the contrary, ours shows clear alignment to the motion conditioning and reliable faithfulness to the first frame. For the Frame Out comparison, we can see that DragAnything [70] might have exaggerated motion when the motion conditioning outside the first frame region is not provided (case 1). Image Conductor [33] cannot faithfully generate the videos. ToRA [87] does not provide a stable Frame Out effect we want.
Figure 8: Extra Frame In Comparisons.
Figure 9: Extra Frame Out Comparisons.
@@ Appendix F Full Canvas Generation Extension
Figure 10: Full Canvas Generation Extension. We show two examples from the teaser image of how our unbounded canvas generation outside the first frame region. The green bounding box is the final generated video that will be cropped in the inference.
The full filed loss implementation makes the generative objective closer to the outpainting model design; however, compared to video outpainting works like [61, 10] which needs full sequence video as inputs and converges to a video-to-video generation task, ours only provides the first frame and focuses on the balance of motion identity-preserving conditioning. Thus, we believe that it is appropriate to name our conditioning as unbounded canvas, instead of the vanilla outpainting. In Fig. 10, we provide the visual content outside the first frame region for two teaser image examples. This is the byproduct of our full field loss implementation, which jointly generates the full canvas region compared with the ground-truth latent in the training. We observe that some examples will show unwanted color distortion for the region outside the first frame region. Since our area of interest is always the first frame region, where we cropped with only the first frame region in the inference. We leave this problem to future works.
@@ Appendix G Limitation
Our method shows promising results, but there are nevertheless some limitations that are worth sharing as shown in Fig. 11. Mainly, this lies in the 3D ambiguity when there is only one point for the motion trajectory. We have to restate that most current works on motion control are single trajectory point-based. One trajectory point for breaking new ID reference information is hard to present the pose ambiguity (see Fig. 11 (c)). Sometimes we want to generate the back view, but we may see the side view cases. Further, one point trajectory is hard to control the size of the ID reference, where sometimes it might be bigger (see Fig. 11 (d) or smaller (see Fig. 11 (e)) than expected. Further, the camera motion lies in the pre-trained model dataset [77] and our filtering method from CUT3R [65] cannot completely remove all videos with camera motion. This leads our model to generate videos with some unwanted camera motion (see Fig. 11 (b)). These issues might be solved by introducing a more robust 3D control system, like camera control or size control for the ID reference.

### v2

@@ Appendix C Additional Experimental Details
In the motion control architecture, one-pixel coordinate as motion images form is insensitive for the VAE encoder to effectively encode and decode; thus, similar to previous works wang2024motionctrl ; wang2024language , we enlarge pixels to a square rectangular box and then execute a 2D dilation algorithm from deblurring domain to increase the perceptibility. The rectangular box border length is 6 pixels for a video with an original height of 384, which means that the size will be scaled with the exact height. Further, to accommodate sparse motion points (1-2 points per object) provided by the user side during the inference, we randomly drop points during the training. The number of tracking points is randomly dropped with the probability of 33% and 60% for each stage respectively. Different from optical flow-based motion control works wang2024motionctrl ; wang2024language ; zhang2024tora , we do not need a padding frame on the last frame to align the total number of frames, where all of our motion conditioning frames represent a spatiotemporal pixel-aligned motion status.
Our inference pipeline can be found in Fig. 5. The model will start from pure noise with the size of the canvas region. The first frame will be expanded to the canvas size by the $\mathbf{B}_{\text{canvas}}$ setting for the top-left and bottom-right expansion pixel number. The motion is conditioned on the canvas size as a whole after converting from the raw pixel spatiotemporal coordinates. Similarly, the ID reference images will be scaled and padded to the same size as the canvas size. Similar to training, after forwarding the video DiT in each timestep, the text tokens and ID tokens will be discarded, but the video DiT in the next timestep will be conditioned on fresh text and ID tokens again. The denoised latent will be decoded by the pre-trained VAE decoder, and then we cropped the pixel space videos to be the exact same shape as the first frame based on $\mathbf{B}_{\text{canvas}}$ setting.
@@
[…]
Works like ReVideo mou2024revideo and VideoAnyDoor tu2025videoanydoor and other similar works zhuang2025get ; bian2025videopainter are conditioned on the full video sequence as inputs, not just the first frame, which means that they are video-to-video generation instead of image-to-video generation we target at. Furthermore, these works are more focused on the insertion and editing in the middle instead of entering and leaving effects we pursue. Thus, we believe that this direction is not an ideal choice for Frame In and Frame Out baselines. An additional condition comparison table can be found in Tab. 5.
@@
FID heusel2017gans measures the distance between the distribution of generated and real frame features, extracted using an Inception network. It is sensitive to both image quality and diversity, and lower scores indicate better generation.
FVD unterthiner2019fvd extends FID to the temporal domain, computing the Fréchet distance between distributions of video-level features extracted using a pre-trained I3D model. It captures spatiotemporal coherence in generated videos and is similarly lower-is-better.
LPIPS zhang2018perceptual quantifies the perceptual similarity between individual frames and their references using deep feature distances from a pre-trained network. Unlike pixel-wise metrics, LPIPS correlates well with human judgment of visual similarity.
The prompt we used for motion description for SkyReels-A2 fei2025skyreels and Phantom liu2025phantom is: Please describe the video in 50 words. Describe the motion of the object clearly and in details, but in the natural and direct language. It is expected that an object will enter and/or exit the image. Describe how the character is moved in and exit, like from the top, left, right, bottom.
Table 6: Human Study with Baseline Methods on Win Rate.
[…]
DragAnything wu2024draganything 83.3%
ImageConductor li2025image 100.0%
ToRA zhang2024tora 73.3%
Phantom liu2025phantom 91.1%
SkyReels-A2 fei2025skyreels 92.2%
Human Study. Further, we conducted a small-scale user study as shown in Tab. 6. We randomly selected 30 videos from our evaluation benchmark set and asked three anonymous human raters to perform pairwise comparisons between our model and baseline methods. The instruction we gave is: You are be shown two generated videos from different models. You need to choose one that appears clearer and better Frame Out and Frame In effect. Meanwhile, we also provide the GroundTruth Motion and BoundingBox information. For the FrameIn cases, you will be given with extra ID reference condition. Frame Out means that an object exists in the first frame condition and is leaving the scene. Frame In means that an identity reference that does not first exist in the first frame, appears in the scene naturally.
In each comparison, the preferred video that achieves a better Frame In and Frame Out effect received a score of 1; the other, 0. A total of 450 comparisons were collected. We report the win ratio, defined as the number of wins divided by the number of comparisons, as a measure of user preference. Higher values indicate a stronger preference for our model. The results align well with the automatic metrics reported in the paper, such as trajectory error and VLM evaluation. The consistency across multiple metrics also provides supporting evidence for the statistical significance of our gains over baselines.
@@ Appendix D Ablation Study
Table 7: Ablation Study on Frame In Comparisons.
@@
As shown in Tab. 7, we compare several different purpose-trained models under different settings. Due to the computation limitations, we train all models with 12K iterations of a batch size of 8. The training is done on stage 2, and all utilize the same pre-trained stage 1 weight (including the baseline). We compare all models on the Frame In evaluation, which is the most representative task for all conditions to be considered.
@@
Full Field Loss Influence. We also provide a model that is not trained with full field loss. This means that the ground-truth target latent provided only has the encoded information from the first frame region instead of the canvas region in the baseline. The region outside the first frame is padded with zero. It is worth noting that the comparison of FID, FVD, LPIPS, Relative DINO, and VLM is only for the first frame region, which means that the padded zero is not included in the evaluation. As we can see from the third row of the Tab. 7, almost all metrics dropped. Here, trajectory error, video segmentation MAE, and relative DINO dropped evidently. We believe that the introduction of full field loss is significant to the final visual generated quality and motion consistency.
Absolute Position Encoding Influence. In our baseline model, we train the model with resized absolute position encoding by trilinear interpolation. Despite this method, another solution to embrace different resolution inputs for both training and inference could be to create a new position encoding each time, where the position encoding for the identity reference is also refreshed each time. However, as shown in the fourth row of Tab. 7, this will lead to a performance drop in all metrics, especially FID and relative DINO. We believe that reusing the learned position encoding by applying a simple trilinear interpolation is helpful for the versatile video resolution inputs, which maintain similar data distribution from the learned fixed position embeddings at the minimum cost.
@@ Appendix E Additional Visualization
Figure 7: More Generated Examples for Our Method.
A more complete sequence of the generated videos and all conditions of the teaser can be found in Fig. 6. Further, we provide more of our generated video samples in Fig. 7. Some demo images or ID references are chosen or cropped from Davis Dataset perazzi2016benchmark and online images. We show multiple different kinds of combinations of Frame In and Frame Out. For the physical interaction, we mean the interaction caused by breaking new ID references to the object that already exists in the first frame. We set all the height and width to be a multiplier of 32 due to the VAE limitation. As shown in Fig. 6 and Fig. 7, the setting for the height and width of the canvas and first frame region can be versatile in aspect ratio and size while generating high fidelity and stable results. Further, we empirically found the generation is stable as long as the canvas height is less than 480 pixels and the canvas width is less than 720 pixels, which is the training resolution for our pre-trained base model, CogVideoX yang2024cogvideox .
As shown in Fig. 8 and Fig. 9, we also present extra qualitative comparisons with the baseline methods. For the Frame In comparison, we can see that baselines like Phantom liu2025phantom and SkyReels-A2 fei2025skyreels cannot understand the Frame In intention we want. The identity reference either already existed in the first frame (cases 2 and 3) or never came to the scene (case 1). Further, the first frame condition is not faithfully considered as the main condition. The scene they generate is similar in elements but different in objects, 3D position, and physical relationship. On the contrary, ours shows clear alignment with the motion conditioning and reliable faithfulness to the first frame. For the Frame Out comparison, we can see that DragAnything wu2024draganything might have exaggerated motion when the motion conditioning outside the first frame region is not provided (case 1). Image Conductor li2025image cannot faithfully generate the videos. ToRA zhang2024tora does not provide the stable Frame Out effect we want.
Figure 8: Extra Frame In Comparisons with Baseline Methods.
Figure 9: Extra Frame Out Comparisons with Baseline Methods.
@@ Appendix F Full Canvas Generation Extension
Figure 10: Full Canvas Generation Showcase. We pick two examples from the teaser image to show the generated results of our unbounded canvas generation outside the first frame region. The green bounding box is the eventual generated video that will be left in the inference.
The full field loss implementation makes the generative objective closer to the outpainting model design; however, compared to video outpainting works like wang2024your ; chen2024follow , which needs full sequence video as inputs and converges to a video-to-video generation task, ours only provides the first frame and focuses on the balance of motion identity-preserving conditioning. Thus, we believe that it is appropriate to name our conditioning as unbounded canvas, instead of the vanilla outpainting. In Fig. 10, we provide the visual content outside the first frame region for two teaser image examples. This is the byproduct of our full field loss implementation, which jointly generates the full canvas region compared with the ground-truth latent in the training. We observe that some examples will show unwanted color distortion for the region outside the first frame region. Since our area of interest is always the first frame region, we cropped with only the first frame region in the inference. We leave this problem to future works.
@@ Appendix G Limitation
Our method shows promising results, but there are nevertheless some limitations that are worth sharing, as shown in Fig. 11. Mainly, this lies in the 3D ambiguity when there is only one point for the motion trajectory. We have to restate that most current works on motion control are single trajectory point-based. One trajectory point for breaking new ID reference information is hard to present the pose ambiguity (see Fig. 11 (c)). Sometimes we want to generate the back view, but we may see the side view cases. Further, one point trajectory is hard to control the size of the ID reference, where sometimes it might be bigger (see Fig. 11 (d) or smaller (see Fig. 11 (e)) than expected. Further, the camera motion lies in the pre-trained model dataset yang2024cogvideox , and our filtering method from CUT3R wang2025continuous cannot completely remove all videos with camera motion. This leads our model to generate videos with some unwanted camera motion (see Fig. 11 (b)). These issues might be solved by introducing a more robust 3D control system, like camera control or size control for the ID reference.